OpenStudy (yureiotaku):

Is this correct? Find the slope intercept form using two points (3,2) and (-2,-3) 2-(-3)/3-(-2)=1=m so slope m=1 ​y=mx+b ​So the slope intercept form is y=1x+(-1) or y=1x-1

OpenStudy (3mar):

May I help?

OpenStudy (yureiotaku):

Of course

OpenStudy (3mar):

That is with my pleasure! I think this would help you.

OpenStudy (yureiotaku):

@3mar Is my slope wrong, I understand to get the slope you have to use the equation y-y1/x-x1. I believe that my slope is correct as I have chosen (-2,-3) as y1 and x1. All I am looking for is a simple confirmation if I have gotten the answer correct or if I didn't to get an explanation on what I am doing wrong.

OpenStudy (mww):

Use the formula described. This is called the point-gradient formula. You just need to know a point on the line and the gradient of that line. The gradient was found in this case through the two pts they gave you. Sub this gradient in to the equation along with ONE of the points (pick whatever is easiest to you.often smaller numbers, but both work) \[y - y_1 = m(x-x_1)\] Use (3,2) say sub in x = 3 and y = 2, with m = 1 What do you get?

OpenStudy (3mar):

No your slope is correct m=1

OpenStudy (3mar):

\[\Huge\color{blue}\checkmark\] Congratulations, @YureiOtaku! I think you got it!

OpenStudy (yureiotaku):

@3mar Thank you, I really just needed confirmation that I was doing the confirmation correctly as this helps me understand if I am getting the slope intercept form correct.

OpenStudy (mww):

the intercept is actually +1. just be careful y - 3 = 1(x-2) so y = x - 2 + 3 = x +1

OpenStudy (3mar):

You hit the right target! \[\Huge\color{red}ツ\]

OpenStudy (3mar):

Time for pray - 35 min - Salam!

OpenStudy (mrnood):

@mww You can see from point (3,2) that y = x-1 the instercept IS -1 y - -3 = x- -2 y + 3 = x+2 y= x-1

OpenStudy (skullpatrol):

$$\huge y=x-1$$ is correct.

OpenStudy (mww):

yes I accidentally mixed the coordinates. silly me.

OpenStudy (phi):

****​So the slope intercept form is y=1x+(-1) or y=1x-1**** Yes, you did it correctly. I assume you know you can also do the calculation for slope by switching the points: (-3-2)/(-2-3) that becomes -5/-5 = 1 (the same number as before) also, we usually don't bother to write 1x, and simply write x. Thus y= x-1 is how we would write the equation.